### v3

# What MLLMs Learn about When they Learn about Multimodal Reasoning: Perception, Reasoning, or their Integration?

Abstract
Evaluation of multimodal reasoning models is typically reduced to a single accuracy score, implicitly treating reasoning as a unitary capability. We introduce MathLens, a benchmark of textbook-style geometry problems that exposes this assumption by operationally decomposing performance into Perception, Reasoning, and Integration. Each problem is derived from a symbolic specification and accompanied by visual diagrams, text-only variants, multimodal questions, and targeted perceptual probes, enabling controlled measurement of each component. Using this decomposition, we show that common training strategies induce systematically different capability profiles that are invisible under aggregate accuracy. Reinforcement learning primarily improves perceptual grounding and robustness to diagram variation, while textual SFT yields gains through reflective reasoning. In contrast, as perception and reasoning improve, a growing fraction of remaining errors fall outside these components and are categorized as integration. These results suggest that apparent progress in multimodal reasoning reflects shifting balances among subskills rather than uniform advancement, motivating evaluation beyond scalar accuracy.
Large Language Models,Multimodal Reasoning,Mathematical Reasoning,Compositionality
1 Introduction
Recent advances in reasoning with Large Language Models (LLMs) have yielded remarkable progress in challenging domains such as Olympiad-level mathematics ((Mathematical Association of America, 2025)), graduate-level scientific question answering ((Rein et al., 2024)), and multi-step program synthesis ((Austin et al., 2021; Chen et al., 2021)). Motivated by these successes, a natural extension is to adapt similar training paradigms to Multimodal Large Language Models (MLLMs), equipping them with reasoning capabilities over both text and visual inputs. Tasks such as mathematical problem solving ((Lu et al., 2024; Wang et al., 2024b)) and visual puzzles ((Dao & Vu, 2025; Ghosal et al., 2024; Feng et al., 2025)) illustrate the potential of this direction, giving rise to methods that adapt Supervised FineTuning (SFT) ((Sun et al., 2025; Chung et al., 2025)) and Reinforcement Learning (RL) ((Deng et al., 2025; Meng et al., 2025)) for multimodal reasoning, including sequentially combining both stages for enhanced reasoning.
Figure 1: MathLens decomposes multimodal reasoning errors into perception, reasoning, and integration, revealing capacity-specific shifts after fine-tuning that are hidden by aggregate accuracy. Each training strategy affects capacities differently; e.g., text SFT yields a minor gain ($\nearrow$) in reasoning but harms ($\downarrow$) integration (model details in Appendix˜D).
However, unlike LLMs where reasoning-oriented training yields consistent gains, multimodal reasoning training exhibits highly variable outcomes. This motivates analysis of how different training strategies influence specific skills. To this end, multimodal reasoning can be analyzed in terms of perception, reasoning, and their integration, each corresponding to a principal source of error. Existing benchmarks, however, rarely adopt such a decomposition and instead primarily report aggregate accuracy, which obscures the distinctions. Some vary input modalities to approximate skill-specific testing, but without strict controls they fail to isolate capacities and offer limited diagnostic value (see Appendix˜B). Consequently, it remains unclear which training strategies benefit multimodal reasoning and why.
A central challenge in addressing this question is the lack of domains where such a decomposition can be carried out with explicit semantic control. Geometry provides a natural setting for this analysis: its symbolic structure makes it possible to precisely specify perceptual facts, reasoning operators, and their interaction, and it already underpins most existing multimodal reasoning benchmarks (Lu et al., 2024; Wang et al., 2024b; Zhang et al., 2024). This level of control is essential for analyzing how training strategies differentially affect perception, reasoning, and their interaction. However, existing benchmarks do not fully exploit this structure to provide controlled, component-wise evaluation.
To close this gap, we present MathLens, a controlled benchmark of geometry problems that separately probes perception, reasoning, and their integration (Figure˜1). Starting from the symbolic semantic state of (Zhang et al., 2023), we build four aligned annotations (Section˜2.2): (i) diagrams rendered from geometric constraints to primarily test perception, (ii) a definitionally equivalent textual description to test reasoning under trivialized perception, (iii) multimodal questions requiring both modalities, and (iv) fine-grained probes targeting recovery of visual details. To further test robustness against visual modifications, MathLens introduces semantic diagram modifications that alter visual form while preserving task correctness. By grounding the benchmark in geometry, MathLens retains authentic task complexity and enables rigorous comparison to prior benchmarks.
MathLens demonstrates that training strategies shape multimodal reasoning capacities in distinct ways. 1) Perception is primarily boosted by reinforcement learning, with larger gains when strong textual SFT has already established reasoning competence (Section˜3.2), while textual SFT itself, despite lacking visual input, indirectly strengthens perception through reflective reasoning (Section˜3.3). 2) Reasoning improves in tandem with perception under RL, but does not exhibit distinct additional gains beyond those coupled improvements (Section˜3.4). 3) Integration remains the least improved capacity: RL offers little benefit, and as perception and reasoning improve, errors increasingly fall outside the scope of the measured perception and reasoning components, and are therefore categorized as integration. (Section˜3.4). 4) Robustness diverges across strategies, with RL enhancing consistency under diagram variation, whereas multimodal SFT reduces robustness through overfitting (Section˜3.5).
Our contributions:
  • Framework for multi-axial evaluation of multimodal reasoning: A new benchmark that separately probes perception, reasoning, and integration, enabling analysis beyond aggregate accuracy.
  • Findings from controlled analysis: Discoveries about how different training objectives and data settings influence the capabilities of Multimodal Language Models.
  • MathLens dataset: A curated benchmark with problems, annotations, and perturbation generators.
[…]
2 MathLens
Most multimodal reasoning benchmarks report only aggregate accuracy, obscuring whether errors stem from perception (extracting information from inputs), reasoning (operating on extracted information), or their interaction. Our benchmark is designed to separately probe these subskills while preserving realistic problem contexts. MathLens comprises 926 geometry problems, each presented with eight visual modifications and an average of $\sim 7.03$ visual probes per problem.
2.1 Dataset formalization
Definitions.
For a problem instance $k$, we consider two latent generative variables: the context semantics $S_{k}$ and the query operator $\varphi_{k}$. The semantics decomposes atomically,
$S_{k}=\{s_{k,1},\ldots,s_{k,m}\},$
where each $s_{k,i}$ encodes a basic fact. For example, in a geometry problem $k$, $s_{k,1}$ could represent ($\angle ABC=50^{\circ}$). Both visual diagram $C^{img}_{k}$ and textual description $C^{txt}_{k}$ context are generated conditionally on $S_{k}$,
$C^{\mathrm{img}}_{k}\sim p(C^{\mathrm{img}}\mid S_{k}),\qquad C^{\mathrm{txt}}_{k}\sim p(C^{\mathrm{txt}}\mid S_{k}).$
The surface question $Q_{k}$ is generated from $\varphi_{k}$ and a subset of atoms,
$Q_{k}\sim p\bigl(Q\mid\varphi_{k},\,S_{k}^{\text{q}}\bigr),\qquad S_{k}^{\text{q}}\subseteq S_{k},$
with $\varphi_{k}=$ (compute $\angle A$) and $S_{k}^{\text{q}}=\{s_{k,1}\}$, the question can be $\angle ABC=50^{\circ}$, what is $\angle A$?" The ground-truth answer is defined by applying the operator to the full semantic state,
$A_{k}=f(\varphi_{k},S_{k}).$
Further, we generate a perception probe set $Q^{\mathrm{perc}}_{k}$ with semantic atoms $s_{k,i}$. Let
$Q^{\mathrm{perc}}_{k}=\{q^{\mathrm{perc}}_{k,i}\}_{i=1}^{m},\qquad a^{\mathrm{perc}}_{k,i}=\llbracket s_{k,i}\rrbracket,$
so each probe targets a single atom of $S_{k}$ and its gold answer is the atom’s valuation. For example, $q^{\mathrm{perc}}_{k,1}=$ “Which angle has measure of 50? A. ABC B. ABD" with $a^{\mathrm{perc}}_{k,1}=$ A. These probes directly test whether a model has recovered the components of $S_{k}$ from the observed context.
Isolating subskills with $S_{k}$ access.
Having access to the semantic state $S_{k}$ allows us to design tests that separately probe perception, reasoning, and their integration (Figure˜2).
Perception. Probe questions $Q^{\mathrm{perc}}k$ with gold answers $a^{\mathrm{perc}}{k,i}$ are derived from the atomic facts listed in $S_{k}$. These probes check if the model can recover the specific facts in $S_{k}$ that are needed to solve the problem from the given input. Errors here indicate perceptual failures.
Reasoning. Given $S_{k}$, we render a textual description $C^{\mathrm{txt}}_{k}$ that directly encodes the relevant details. Evaluating on $(C^{\mathrm{txt}}_{k},Q_{k})$ trivializes perception, so the task reduces to applying $\varphi_{k}$ correctly. Errors here isolate reasoning competence, free from perceptual confounds.
Integration. We isolate integration effects by combining accuracy on $(C^{\mathrm{img}}_{k},Q_{k})$ with auxiliary perception and reasoning measures. Conditional on success in perception probes and text-only reasoning, any remaining errors on the full task are treated as integration errors; this defines integration as a residual category after controlling for the measured perception and reasoning components under standard, non-exhaustive coverage assumptions.
Accordingly, integration should be interpreted as a conditional residual rather than a fully disentangled construct: it captures failures not explained by the measured perception probes and text-only reasoning, but may still subsume unmeasured perceptual factors.
Additional uses of $S_{k}$.
First, $S_{k}$ allows construction of questions that require context. We define $Q^{\prime}_{k}$ by restricting its atoms to exclude those in the context latent $S_{k}^{\text{c}}$:
$Q^{\prime}_{k}\sim p(Q\mid\varphi_{k},S_{k}^{\text{q}^{\prime}}),\qquad S_{k}^{\text{q}^{\prime}}\subseteq S_{k},\qquad S_{k}^{\text{q}^{\prime}}\cap S_{k}^{\text{c}}=\varnothing.$
By design, $Q^{\prime}_{k}$ cannot be answered from the query alone and thus forces reliance on $C_{k}$.
Second, the symbolic representation $S_{k}$ allows us to apply semantic perturbations of the context. Let $\tau\in\mathcal{T}_{\mathrm{AP}}$ denote a transformation from the set of admissible perturbations (e.g., relabeling points or rotating a diagram). Applying $\tau$ yields a perturbed specification $S^{\prime}_{k}=\tau(S_{k})$, while preserving the problem semantics so that
$f(\varphi_{k},S_{k})=f(\varphi_{k},S^{\prime}_{k}).$
Re-rendered contexts ($C^{\mathrm{img}}_{k^{\prime}}\sim p(C^{\mathrm{img}}\mid S^{\prime}_{k})$ and $C^{\mathrm{txt}}_{k^{\prime}}\sim p(C^{\mathrm{txt}}\mid S^{\prime}_{k})$) serve as systematic distractors. We measure robustness by prediction agreement across contexts from $S_{k}$ and $S^{\prime}_{k}$, testing semantic invariance beyond pixel-level augmentation. Pixel-level perturbations are insufficient, since they often cause prediction shifts driven by abnormal appearance rather than semantic change.
Figure 3: Sample data generation process in MathLens. From a semantic state representation, we build controlled text descriptions, perception probes, and questions with no overlap with visual content. Also, new diagrams are rendered from the semantic state to avoid visual familiarity effects [^2] Here, an obtuse angle is drawn acute to prevent visual estimation and enforce geometric deduction, consistent with Olympiad practice (e.g., “not drawn to scale” in AMC)..
2.2 Data generation pipeline
Thus, we build on latent semantics $S_{k}$ for a set of practical multimodal reasoning problems, matching problem domain and complexity with popular benchmarks in the literature ((Zhang et al., 2024; Lu et al., 2024)). Each instance is first specified symbolically as $S_{k}$ and $\varphi_{k}$, then rendered into its observable forms: diagrams $C^{\mathrm{img}}_{k}$, textual descriptions $C^{\mathrm{txt}}_{k}$, and multiple types of questions.
Data source (Figure˜4 a).
We build on FormalGeo-7K ((Zhang et al., 2023)), which provides symbolic annotations for geometry problems. Each diagram cue or condition is encoded as a predicate (e.g., $\textsc{Collinear}(AB,BC)$, $\textsc{EqualLength}(AB,CD)$), forming the semantic state $S_{k}$. This yields realistic problems with formal representations from which we build the required artifacts.
Questions (Figure˜4 b).
We represent each question $Q_{k}$ as clauses $S_{k}^{\text{q}};\varphi_{k}]$, with facts $S_{k}^{\text{q}}$ and goal operator $\varphi_{k}$. For a strictly multimodal question, we drop clauses overlapping with the context $S_{k}^{\text{c}}$:
$Q^{\prime}_{k}=f_{\text{q}}\bigl([(S_{k}^{\text{q}}\setminus S_{k}^{\text{c}});\varphi_{k}]\bigr),$
where $f_{\text{q}}$ linearizes them into natural language. Thus $Q^{\prime}_{k}$ requires contextual information to solve.
Textual descriptions (Figure˜4 c).
Each detail $s_{k,i}\in S_{k}$ is mapped by template function $f_{\text{d}}$ to clause,
$C^{\mathrm{txt}}_{k}=\text{concat}_{i}f_{\text{d}}(s_{k,i}),$
yielding a faithful textual rendering of formal representation $S_{k}$ without stylistic variation (e.g., $\textsc{EqualLength}(AB,CD)\!\to$ “Segment $AB$ is equal in length to segment CDCD”).
Perception probes (Figure˜4 d).
Each atomic detail $s_{k,i}\!\in\!S_{k}$ is converted to a probe via templating function $f_{\text{p}}$ with gold answer $\llbracket s_{k,i}\rrbracket$:
$Q^{\mathrm{perc}}_{k}=\{f_{\text{p}}(s_{k,i})\}_{i},\qquad a^{\mathrm{perc}}_{k,i}=\llbracket s_{k,i}\rrbracket.$
For example, $\textsc{Collinear}(AB,BC)$ yields “Which points are collinear? A. ABC B. ABD” with answer A and negative B. Thus probes directly test recovery of $S_{k}$ from context.
Diagram rendering (Figure˜4 e).
Each clause $s_{k,i}\!\in\!S_{k}$ is converted into geometric constraints (e.g., $\textsc{Perpendicular}(AB,BC)$ $\to(x_{a}-x_{b})(x_{c}-x_{b})+(y_{a}-y_{b})(y_{c}-y_{b})=0$). A numerical solver computes coordinates that satisfy all constraints, which are then rendered into diagrams. We then manually filter outputs to remove artifacts such as occlusions or overlaps.
Diagram modifications.
To test robustness to distractions, we generate diagram variants by altering the semantic state $S_{k}$, including adding auxiliary elements, applying flips or rotations, merging instances, and relabeling points, while preserving the ground-truth answer. All variants are manually screened for visual validity.
2.3 MathLens-NS: A non-symbolic set
The main dataset, MathLens, focuses on geometry problems, a domain selected intentionally for its well-defined symbolic structure and precise experimental controllability. This choice is methodological: geometry provides a natural and practical setting for isolating and analyzing the multimodal reasoning mechanisms studied in this work.
In addition, we construct MathLens-NS, a supplementary collection of multimodal reasoning problems drawn from diverse domains. Relative to MathLens, this dataset necessarily trades formal semantic rigor for broader domain inclusion, reflecting the intrinsic differences between geometry and less structured multimodal reasoning settings. As a result, MathLens-NS does not admit explicit semantic state specifications; instead, it is curated from prior sources through a consistent and rigorous filtering pipeline to ensure reliability and diversity. Curation procedures and experimental results are provided in Appendices˜C and E.
Figure 4: Impact of multimodal reasoning training on MathLens. We evaluate pretrained backbones alongside models finetuned for multimodal reasoning, reporting accuracy (%). MathLens is sensitive to gains from multimodal reasoning-oriented finetuning.
3 Experiments
We study open-weight multimodal reasoning models in the 7-9B range and their backbone counterparts, motivated by the transparency of open models and the fact that most released multimodal reasoners fall within this scale. We evaluate seven model families, totaling 13 checkpoints, including both SFT-only and RL-finetuned variants where available. We include Qwen-2.5-VL (Bai et al., 2025) and GLM-4.1V-Base (Hong et al., 2025) as backbone models. Direct multimodal Reinforcement Learning (RL) models comprise VL-Rethinker (Wang et al., 2025) and ShareVL-R1 (Yao et al., 2025). Models trained with multimodal Supervised Finetuning (SFT) followed by multimodal RL include Vision-R1 (Huang et al., 2025) and R1-Onevision (Yang et al., 2025), while those using textual SFT followed by multimodal RL include Revisual-R1 (Chen et al., 2025) [^3] Revisual-R1 adds additional textual RL, yet we group it as textual SFT $\rightarrow$ multimodal RL for consistency. and Open-Vision-Reasoner (OVR) (Wei et al., 2025). Finally, GLM-4.1V-Thinking (Hong et al., 2025) falls outside these categories because its training data is undisclosed.
We additionally evaluate six larger models, yielding eight runs in total by including both Gemini-2.5-Flash (Comanici et al., 2025) and Claude 4 Sonnet (Anthropic, 2025) in “thinking” and “non-thinking” modes: Qwen-2.5-VL-72B, VL-Rethinker-72B, Gemini-2.5-Flash, Claude 4 Sonnet, and GPT-O3/4O (OpenAI, 2025). For each model, we follow the recommended decoding configurations to generate both reasoning and final answers. Most models employ greedy decoding and yield deterministic outputs. Full experimental details and results are in Section˜E.3.
Figure 5: Correlation of MathLens with popular benchmarks. MathLens shows high correlation with standard multimodal reasoning benchmarks.
Weak textual reasoning in existing multimodal SFT models.
We take it as baseline that current multimodal SFT models show limited reasoning following (Chen et al., 2025): their data are easier than text-only sets, leading to weaker performance on reasoning benchmarks. Building stronger data is hindered by the lack of open multimodal reasoning traces (see Section˜E.1 for discussion).
Statistical significance.
Because our primary claims concern changes induced by finetuning, we report paired bootstrap confidence intervals for all pre-post finetuning deltas over problem instances (Figure˜6, 7, and 10). Unless otherwise noted, all evaluations use deterministic decoding; accordingly, reported uncertainty reflects benchmark sampling variability rather than decoding stochasticity. Raw accuracies and diagnostic visualizations (Figure˜4, 9, and 13) are reported descriptively and are not subjected to statistical inference.
3.1 Validating MathLens
Before analyzing finetuning effects, we validate MathLens as a multimodal reasoning benchmark by showing it captures the same skills as established benchmarks.
Sensitivity to finetuning.
We assess whether MathLens reflects gains from multimodal reasoning finetuning by comparing pretrained backbones with their finetuned variants on its downstream task of solving geometry problems from diagrams. As shown in Figure 4, finetuned models consistently surpass backbones, demonstrating that MathLens is sensitive to these adaptations.
Correlation with established benchmarks.
To test whether MathLens captures patterns consistent with prior benchmarks, we compare it to MathVista ((Lu et al., 2024)), MathVerse ((Zhang et al., 2024)), and MathVision ((Wang et al., 2024b)) by correlating model accuracies across eight models (Figure 5, left; full results in Appendix). MathLens shows strong Spearman’s $\rho$ correlation with MathVista ($\rho=0.83$) and MathVerse ($\rho=0.86$), confirming alignment with established benchmarks, while correlation with MathVision is weaker but still positive ($\rho=0.67$). Importantly, our goal is not a new downstream benchmark but a controlled, decomposition-focused resource; high correlations therefore underscore its consistency.
3.2 How much of the gain is explained by improvement in textual reasoning?
Figure 6: Performance gains by input modality. Bars show percentage point shifts from finetuning for text versus diagram inputs. Visual gains exceed textual ones when models are primed with strong reasoning (textual SFT).
To isolate the role of textual reasoning, we compare performance on textual descriptions $C^{\mathrm{txt}}_{k}$ and visual diagrams $C^{\mathrm{img}}_{k}$, which encode identical information conditioned on the question $Q_{k}$. This parallel annotation lets us track how multimodal training affects each modality. We report accuracy differences (percentage points) before and after multimodal finetuning.
Figure˜6 shows that textual SFT mainly improves textual reasoning, while multimodal RL applied afterward yields larger multimodal gains by shifting its effect to perception and integration. Direct multimodal RL without textual SFT gives only modest improvements in both modalities, and poor-quality multimodal SFT degrades multimodal performance more than textual reasoning, with RL only partially recovering the gap.
Finding 1: Multimodal RL impact varies with textual reasoning strength. With strong textual SFT it mainly boosts perception; without it, it modestly improves both modalities (Figure˜6).
3.3 How much of the gain is explained by improvement in perception?
Figure 7: Perception performance shifts from finetuning in percentage points. Except for multimodal SFT, all methods contribute to increased perception capacity.
Figure 8: Sample distribution over reflection count. Reflection is more frequent in false→true cases than in true→true cases, showing that perception gains from textual SFT arise partially from reflective reasoning.
Next, we evaluate how multimodal reasoning training affects low-level perception. For each problem $Q_{k}$, we use the perception probes $Q^{\mathrm{perc}}_{k}={q^{\mathrm{perc}}_{k,1},\ldots,q^{\mathrm{perc}}_{k,j}}$. Figure˜7 shows that all RL models contribute to better perception required for geometry problem solving. Hence, we conclude that perception is elicited even by training with correctness reward signals on the downstream problems.
Figure 9: Robustness to semantic-level visual modifications. Box plots show accuracy on modified diagrams. Points report accuracy on the unmodified base diagrams and the overall consistency score. Multimodal RL improves consistency, whereas multimodal SFT reduces it.
Figure 10: Finetuning shifts by diagram type. Bars show percentage point changes for existing versus OOD rendered diagrams.
How textual SFT improves perception.
Textual SFT does not use visual inputs, yet it enhances perceptual performance. These gains suggest influences beyond direct perceptual learning. We hypothesize that one contributing factor is that enhanced reasoning alters how the model interprets ambiguous visual evidence. In particular, stronger reasoning promotes cognitive strategies such as reflection and self-correction, which allow the model to revise initial perceptual judgments.
[…]
Finding 2: Textual SFT also improves perception partially by fostering reflective reasoning (Figure˜8).
3.4 Error type analysis
MathLens’s annotation designs enable systematic categorization of model outputs into the following error categories: (1) Perception & Reasoning, failures on both perception probes and textual reasoning; (2) Perception, failure on perception probes but correct reasoning from text; (3) Reasoning, correct perception probes but failed text-based reasoning; (4) Integration, correct perception and text in isolation but failure on the combined multimodal task. Correct categories: (5) Trivial, solvable from text alone; and (6) Rest, all other correct cases.
Figure˜1 (see full results in Figure˜14 of Appendix) shows that RL primarily reduces perception and reasoning errors on the same problems, indicating correlated gains. Yet many of the reductions manifest as Integration errors, suggesting that finetuning often shifts them into failures attributed to coordination.
Finding 3: RL improves perception and reasoning in a correlated manner, and causes residual errors to be increasingly categorized as integration once the other components improve (Figure˜1).
3.5 Does multi-modal finetuning effect model’s robustness to visual inputs?
We extend the visual familiarity analysis from Section˜3.1 by testing robustness under controlled diagram variations. Instead of pixel-level augmentations (e.g., blurring), we apply semantic modifications to the geometric specification, isolating structural variation without introducing low-level artifacts (see Appendix˜C).
Downstream accuracy can mask familiarity effects.
Figure˜10 shows that while most training methods are stable under familiar diagrams, multimodal SFT suffers a sharp drop on modified diagrams. This indicates that similar downstream accuracy on public benchmarks can hide reliance on visual familiarity.
RL improves visual consistency.
To quantify robustness, we use Consistency Rate (CR) ((Zhao et al., 2024)), the expected agreement of predictions across perturbations of the same diagram:
$\Delta$ $=\mathbf{1}\!\left[M(Q^{\prime}_{k},\bar{C}_{k,i})=M(Q^{\prime}_{k},\bar{C}_{k,j})\right].$ (1)
$CR$ $=\mathbb{E}_{Q^{\prime}_{k}\sim Q^{\prime}}\Bigg[\mathbb{E}_{\begin{subarray}{c}\bar{C}_{k,i},\,\bar{C}_{k,j}\sim\bar{C}_{k}\\ i\neq j\end{subarray}}\Delta\Bigg].$ (2)
where $\bar{C}_{k}$ is the set of diagram variants for question $Q^{\prime}_{k}$.
Figure˜9 shows that accuracy on base diagrams is generally higher than on modified ones, indicating vulnerability to semantic perturbations. Consistency increases after multimodal RL, suggesting robustness to variation. By contrast, multimodal SFT lowers consistency.
Finding 4: Multimodal RL improves visual consistency under structural variations (Figure˜9).
3.6 Which perception skills benefit from multimodal reasoning training?
We analyze perception probe performance (Section˜3.3) by relation type, using a fine-grained categorization of geometric cues (see Section˜E.3 for full results). This analysis indicates that gains from multimodal reasoning training are not uniform across perceptual skills. Relations such as cocircular, parallel, and collinear, which depend on simple local geometric primitives, show consistent improvement. In contrast, relations such as same_angle, val_angle, triangle, and quadrilateral benefit only when textual supervision is combined with multimodal reinforcement learning, reflecting their reliance on multi-constraint reasoning and symbol–geometry alignment. By comparison, same_length, val_length, and perpendicular remain challenging, likely due to spatially displaced cues or visually ambiguous annotations. Detailed per-category results are reported in the appendix.
Finding 5: Perception gains are uneven. Multimodal reasoning training preferentially improves skills driven by direct geometric cues, while perception tasks that rely on symbolic marks or long-range visual associations show limited gains.
4 Conclusion
We introduced MathLens, a controlled benchmark that disentangles perception, reasoning, and integration in multimodal reasoning. We find that reinforcement learning primarily improves perception, especially when paired with textual supervision, while textual SFT strengthens perception indirectly through reflective reasoning. Reasoning gains largely follow perceptual improvements, leaving integration as a dominant source of residual error, and robustness diverges across strategies: RL improves consistency under diagram variation, whereas multimodal SFT reduces it due to overfitting.
Looking ahead, our results motivate future work on architectures and training strategies that more directly address cross-modal integration and perceptual capacity.
Broader Impact
Our goal is to improve the evaluation and understanding of multimodal reasoning models. By decomposing performance into perception, reasoning, and their integration, our benchmark provides more diagnostic insight than aggregate accuracy and helps clarify which subskills are responsible for observed gains.
Our work focuses on mathematical geometry problems and is constructed entirely from publicly available problem sources and synthetic derivations. It does not involve personal data, sensitive attributes, or application domains with direct societal deployment. As such, the benchmark primarily serves as a scientific analysis tool for the research community.
As with any benchmark, there is a risk that future work may over-optimize for its specific structure rather than broader generalization. To reduce this risk, the benchmark is designed with explicit subskill decompositions and controlled variants intended for diagnostic analysis rather than leaderboard-driven optimization. We do not anticipate additional ethical or societal risks beyond those commonly associated with advances in multimodal machine learning.
Appendix A Related Work
Evaluation of multimodal reasoning capacity.
Multimodal reasoning is a compositional process that requires the integration of perception and abstract reasoning ((Li et al., 2025)). Inspired by textual reasoning benchmarks, recent multimodal reasoning datasets focus on verifiable domains such as mathematics ((Lu et al., 2024; Wang et al., 2024b; Zhang et al., 2024)), scientific diagrams ((Yue et al., 2024a; Hao et al., 2025)), and charts ((Wang et al., 2024c)). However, most of these benchmarks report only a single downstream accuracy per model, without systematic means to identify the source of errors. Some further provide manually annotated error-type analyses ((Zhang et al., 2024; Hao et al., 2025)), but these rely on non-standard category definitions across benchmarks, depend on post-hoc semantic inspection of reasoning traces rather than causal diagnosis, and suffer from annotator variance. In addition, the necessity of multimodal context is often left unverified ((Yue et al., 2024b)). Finally, many benchmarks extract problems from publicly available sources such as textbooks, raising risks of data leakage and familiarity bias. To overcome these limitations, our study builds on FormalGeo-7K ((Zhang et al., 2023)), which provides formal abstractions for both context and goal in practical mathematical geometry problems, enabling rigorous and fine-grained analysis.
Training Methods for Multimodal Reasoning.
Approaches to adapting MLLMs for multimodal reasoning typically fall into three types: multimodal supervised finetuning, where models are trained on paired image–text inputs with ground-truth reasoning traces and answers; multimodal reinforcement learning, where models are optimized with rewards from verifiable outcomes or reasoning-trace feedback; and textual supervised finetuning, where models are tuned on large-scale textual reasoning corpora without multimodal context. These strategies are often applied sequentially or in combination, such as multimodal SFT followed by multimodal RL (e.g., (Huang et al., 2025) and (Yang et al., 2025)) to improve robustness, or textual SFT followed by MM-RL (e.g., (Chen et al., 2025) and  (Wei et al., 2025)) to transfer reasoning priors into multimodal domains. In addition, some models adopt direct RL without prior SFT (e.g., (Deng et al., 2025)(Meng et al., 2025)(Wang et al., 2025), and (Yao et al., 2025)), demonstrating that reinforcement learning alone can yield competitive reasoning performance. Closed-weight systems (e.g., (OpenAI, 2025)(Comanici et al., 2025), and (Anthropic, 2025)) also report strong multimodal reasoning ability, although their training data and pipelines remain undisclosed.
Appendix B Comparison with Existing Benchmarks
Figure 11: Example of MathVerse text descriptions. The textual description fails to fully encode geometric relations, requiring external information to be inferred from the diagram. This incompleteness makes it unsuitable for evaluating pure reasoning capacity.
Figure 12: Inconsistencies in manual error analyses across benchmarks. Pie charts show variation in error categories across multimodal reasoning datasets, underscoring the lack of standardized criteria. Analyses are typically restricted to a single model. Figures are taken from the respective papers (Hao et al., 2025; Wang et al., 2024b; Zhang et al., 2024).
Comparison with MathVerse.
Multimodal mathematical problems are a popular testbed for evaluating multimodal reasoning. Among existing benchmarks ((Lu et al., 2024; Wang et al., 2024b; Yue et al., 2024a)), the closest to ours is MathVerse ((Zhang et al., 2024)), which, like our benchmark, primarily focuses on mathematical geometry problems and partly draws from Geometry3K ((Lu et al., 2021)), a subset of FormalGeo-7K ((Zhang et al., 2023)). MathVerse also provides multiple input modalities ranging from text-only to vision-only, enabling skill-specific evaluation of multimodal reasoning models, which is conceptually aligned with our design.
However, MathVerse does not meet the criteria necessary for rigorous capacity isolation: 1) it relies on curated diagrams, making it vulnerable to data leakage or familiarity effects; 2) its text descriptions are incomplete for evaluating pure reasoning, often requiring external information to be inferred from diagrams (see Figure˜11); 3) overlap between question and diagram content is not explicitly controlled. Due to these limitations, analyses in this paper cannot be reproduced with the same rigor using prior benchmarks.
Benefits of automatic error analysis.
Error type analysis offers actionable insights into model weaknesses and guides directions for improvement. However, in multimodal reasoning research, such analyses lack standardization and are almost always performed manually. Consequently, as shown in Figure˜12, categories and criteria vary across datasets, making it difficult to generalize findings. Moreover, manual analysis is costly and typically applied to a single model, quickly becoming outdated along with its conclusions. By contrast, MathLens’s automatic analysis pipeline enables consistent, scalable, and up-to-date error categorization across models and datasets.
Appendix C Data Generation Details
C.1 MathLens
Diagram rendering.
Each clause $s_{k,i}\!\in\!S_{k}$ is mapped to a corresponding set of algebraic constraints. For instance, $\textsc{Perpendicular}(AB,BC)$ is encoded as ($(x_{a}-x_{b})(x_{c}-x_{b})+(y_{a}-y_{b})(y_{c}-y_{b})=0$). Taken together, the constraints form a nonlinear system defining admissible coordinates for all points. We solve this system using the sequential least-squares programming routine (SLSQP) in scipy.optimize.minimize ((Virtanen et al., 2020)), which searches for coordinates that minimize the residual of all constraints subject to feasibility bounds. To improve robustness, initial coordinates are randomly sampled, and optimization is repeated until convergence. Up to ten attempts with different random seeds are allowed, after which the sample is discarded if no feasible solution is found. Valid solutions are rendered with a matplotlib-based backend ((Hunter, 2007)) that draws points, line segments, arcs, and annotations according to the computed geometry. Since automatic optimization may still yield degenerate layouts (e.g., overlapping vertices, occluded labels, or extreme aspect ratios), we apply a post-processing step in which such outputs are manually filtered to preserve clarity and readability.
Diagram modification.
For each geometry problem, we generate eight diagram variants: (1) the original diagram from the source dataset, (2) a rendered diagram generated directly from the symbolic representation, and six symbolic modifications: (3) add_shapes, which inserts 1–3 random shapes (triangles or quadrilaterals), (4) add_lines, which inserts 1–3 random lines between existing points, (5) flip, which mirrors the canvas while preserving label orientation, (6) rotate, which rotates the canvas while keeping labels upright, (7) merge, which concatenates the current diagram with another randomly chosen one and revises labels accordingly, (8) rename, which replaces the label set (e.g., A,B,C → X,Y,Z). All auto-rendered figures are manually filtered after generation. If any version is invalid or visually unsuitable (e.g., severe occlusions, degenerate angles), the entire problem is discarded.
C.2 MathLens-NS
MathLens-NS is a curated, re-annotated benchmark that extends the scope of MathLens to a broader range of images and problem domains. It includes 107 problems, each paired with an average of $\sim 7.96$ visual probe questions. Representative samples are shown in Figure˜15 and Figure˜16, with the data generation procedure detailed below.
Sample collection.
Problems are sourced from six established multimodal reasoning datasets: BLINK ((Fu et al., 2024)), V* ((Wu & Xie, 2024)), SpatialEval-Real ((Wang et al., 2024a)), MMMU-Pro ((Yue et al., 2024b)), EMMA ((Hao et al., 2025)), and MathVista ((Lu et al., 2024)), with mathematical geometry items excluded using metadata. We retain only multiple-choice problems to maintain consistency. Problems requiring more than two images are discarded, and dual-image inputs are concatenated into single images to ensure compatibility with models that do not support multi-image inputs.
Data filtering.
Problems that appear multimodal may in fact be solvable from text-only correlations ((Yue et al., 2024b)), while others that look complex may reduce to simple pattern matching. To guard against such shortcuts, we use model-based validation to test multimodality and reasoning requirements. Each problem is evaluated with Gemini-2.5-Flash ((Comanici et al., 2025)) under three settings: (1) full input, (2) text-only input without the image, and (3) full input with reasoning disabled. We generate eight responses per setting at temperature 0.6 to capture variability.
A problem passes the multimodality check if text-only accuracy does not exceed chance (1/$k$ for $k$-way choice, e.g., 25% for 4 options). It satisfies the reasoning check if accuracy with reasoning disabled falls below chance. We enforce solvability by discarding problems for which the model fails to answer correctly in all eight full-input attempts. Image-level deduplication is then applied to remove visually similar items.
Finally, human annotators enforce validity by independently solving each filtered problem, retaining only those with clearly determinable correct answers. This process reduces the initial pool of $\sim$2,000 problems to 200 high-quality samples that meet all four criteria.
Manual annotation.
We generate seed annotations with Gemini-2.5-Flash, then manually revise them to remove hallucinations and add missing details needed for solvability. Textual descriptions are structured in a scene-graph format and decomposed into atomic clauses, from which perception probes are automatically derived.
Appendix D Implementation Details and resources
Models.
Details of the model configurations and corresponding sources are provided in Table˜1.
Model Type Source
7–9B Open-Weight
VL-Rethinker backbone Qwen/Qwen2.5-VL-7B-Instruct
$m$-rl TIGER-Lab/VL-Rethinker-7B
ShareVL-R1 backbone Qwen/Qwen2.5-VL-7B-Instruct
$m$-rl HuanjinYao/R1-ShareVL-7B
R1-OneVision backbone Qwen/Qwen2.5-VL-7B-Instruct
$m$-sft Fancy-MLLM/R1-Onevision-7B
$m$-rl Fancy-MLLM/R1-Onevision-7B-RL
Vision-R1 backbone Qwen/Qwen2.5-VL-7B-Instruct
$m$-sft Osilly/Vision-R1-CI-7B
$m$-rl Osilly/Vision-R1-7B
Revisual-R1 backbone Qwen/Qwen2.5-VL-7B-Instruct
text-sft csfufu/Revisual-R1-Coldstart
$m$-rl csfufu/Revisual-R1-final
OVR backbone Qwen/Qwen2.5-VL-7B-Instruct
text-sft Kangheng/OVR-7B-ColdStart
$m$-rl Kangheng/OVR-7B-RL
GLM-4.1V backbone zai-org/GLM-4.1V-9B-Base
$m$-rl zai-org/GLM-4.1V-9B-Thinking
72B Open-Weight
VL-Rethinker backbone Qwen/Qwen2.5-VL-72B-Instruct
$m$-rl TIGER-Lab/VL-Rethinker-72B
Closed-Weight
OpenAI backbone GPT-4O (gpt-4o_2024-11-20)
thinking GPT-O3 (o3_2025-04-16)
Gemini backbone gemini-2.5-flash (thinking=disabled)
thinking gemini-2.5-flash (thinking=enabled)
Claude backbone claude-4-sonnet (thinking=disabled)
thinking claude-4-sonnet (thinking=enabled)
Table 1: Model configurations studied in this work.
Model configuration for Figure˜1.
All models are fine-tuned from Qwen-2.5-VL-7B as the backbone MLLM. VL-Rethinker represents the direct RL setting, Vision-R1 serves as the multimodal SFT model, and Revisual-R1 corresponds to the textual SFT model. Vision-R1 and Revisual-R1 further include their respective RL-extended variants.
Hyperparameters & computation.
We use Eureka ML Insights Framework ((Balachandran et al., 2024)) for reproducible evaluation. We run 7–9B parameter models on four NVIDIA A100 80GB GPUs, and 72B models on eight. Generation is accelerated and parallelized with the vLLM ((Kwon et al., 2023)) library. Most experiments use greedy decoding (temperature $0.0$) for deterministic outputs. For models that otherwise suffer from text degeneration through severe repetition (e.g., (Huang et al., 2025)), we apply stochastic decoding with temperature $0.6$ and top-$p$ $0.65$. The default maximum generation length is 32,768 tokens to accommodate long reasoning chains. For OVR ((Wei et al., 2025)), we extend this limit to 48,000 tokens due to frequent truncations at lower cutoffs.
Visual resources.
Visual icons used in Footnote˜2 are adapted from flaticon.com.
Large Language Model Usage.
LLMs (ChatGPT, GPT-4/5 class and Claude 4 Sonnet) were employed to refine phrasing, improve clarity, and standardize style in sections of the manuscript, but all scientific ideas, experiments, and analyses were conceived, executed, and validated by the authors. LLMs were also used in a limited capacity to assist with literature discovery (e.g., surfacing related work for manual screening). All substantive content decisions, experiment design, and result interpretation remain entirely author-driven.
Appendix E Experiment
E.1 Preliminaries
Weak textual reasoning in multimodal SFT models.
Prior work ((Sun et al., 2025)) highlights that multimodal SFT datasets are considerably easier than textual SFT datasets, which contributes to weaker textual reasoning capacity in trained models. For example, the Vision-R1 dataset ((Huang et al., 2025)) averages 821.5 tokens per reasoning trace with a 96.0% pass rate, whereas the text-only DeepMath dataset ((He et al., 2025)) averages 8,207.8 tokens with a 75.0% pass rate. This difference suggests that multimodal SFT data require substantially less reasoning effort. Consistent with this, multimodal SFT models trained on such data underperform on standard reasoning benchmarks compared to textual SFT models. Finally, while textual SFT data can be constructed by distilling reasoning traces from large language models ((Guo et al., 2025)), no comparably strong multimodal reasoning models with open reasoning traces currently exist, making effective multimodal SFT data generation particularly challenging.
E.2 Further Insights
Insight on visual familiarity effects.
In Figure 3 (left) of the main paper, we additionally evaluate MathLens-E, a variant that uses the same geometry problems but replaces rendered diagrams with existing diagrams from the original sources. Interestingly, MathLens-E correlates strongly with MathVision ($\rho=0.81$), while showing weaker correlations with MathVista and MathVerse. This difference suggests that models may leverage visual familiarity with diagram styles from textbooks or public tests when tackling MathVision, an advantage that disappears with freshly-rendered diagrams. This underscores that high benchmark accuracy does not necessarily indicate strong multimodal reasoning, as performance may be inflated by visual familiarity effects from training data.
E.3 Full Results
Figure 13: Perception probe results by question type. Improvements differ across skills: angle-related and cocircularity tasks improve with multimodal RL, line-length tasks show little change, and polygon detection improves only with textual SFT. A likely factor is whether notations are colocated with the geometric element they describe or spatially separated.
Perception probe results by relation type.
We report perception probe performance decomposed by question type in Figure˜13. Each category groups probes according to the geometric relation being tested (e.g., angle relations, length relations, polygon structure). Performance gains differ substantially across categories. Angle-related relations and cocircularity show consistent improvement under multimodal reinforcement learning, while line-length relations exhibit limited change across training variants. Polygon-level relations, including triangle and quadrilateral detection, improve primarily when textual supervised fine-tuning is included. These differences appear correlated with the visual presentation of the target cue. Relations whose defining annotations are colocated with the relevant geometric element tend to benefit more from multimodal training, whereas relations relying on spatially separated marks or distant comparisons remain challenging.
Model MathVista MathVerse MathVision MathLens-E MathLens
Qwen2.5-VL-7B 68.2 46.3 25.1 34.7 33.2
R1-Onevision-7B 64.1 46.4 29.9 35.4 29.8
$M$-Eureka-Qwen-7B 73.0 50.3 26.9 34.0 31.3
ThinkLite-VL-7B 74.3 52.2 29.9 34.3 32.9
R1-ShareVL-7B 75.4 52.8 29.5 36.7 35.4
VL-Rethinker-7B 74.9 54.2 32.3 37.6 35.7
Qwen2.5-VL-72B 74.8 57.2 38.1 41.6 41.0
VL-Rethinker-72B 80.4 63.5 44.9 47.6 45.0
Table 2: Full results of all models used to produce the correlation plot in Figure˜5.
Correlation plot details.
Table 2 reports the full benchmark scores for all models used in the correlation analysis of Figure˜5. Results for MathVista, MathVerse, and MathVision are drawn from prior work ((Wang et al., 2025)). Consequently, the set of models differs from our main evaluation and includes additional variants such as MM-Eureka ((Meng et al., 2025)) and ThinkLite-VL ((Deng et al., 2025)). These models were excluded from the main analysis for two reasons: (i) to maintain a balanced number of models across categories, particularly those trained with direct RL, and (ii) to focus on the stronger-performing models on other benchmarks.
Model Variant Text Raw Base Add Add Flip Merge Rename Rotate Consistency
lines shapes
VL-Rethinker-7B Backbone 38.9 34.7 33.2 26.1 28.1 30.6 26.9 29.4 31.4 31.0
$M$-RL 41.1 37.6 35.7 31.4 29.4 34.7 30.6 33.5 35.5 37.3
ShareVL-R1-7B Backbone 38.9 34.7 33.2 26.1 28.1 30.6 26.9 29.4 31.4 31.0
$M$-RL 42.1 36.7 35.4 31.1 32.7 33.9 33.7 35.1 35.3 41.3
R1-OneVision-7B Backbone 38.9 34.7 33.2 26.1 28.1 30.6 26.9 29.4 31.4 31.0
$M$-SFT 36.7 35.4 29.8 21.9 26.1 29.7 25.3 27.6 29.2 24.0
$M$-RL 35.9 33.0 32.0 25.1 28.5 34.0 23.3 30.7 32.3 28.5
Vision-R1-7B Backbone 38.9 34.7 33.2 26.1 28.1 30.6 26.9 29.4 31.4 31.0
$M$-SFT 35.2 34.3 26.8 19.1 21.7 27.8 20.6 23.1 28.0 21.1
$M$-RL 48.6 44.5 36.0 33.0 31.5 36.6 30.6 34.2 37.8 39.4
Revisual-R1-7B Backbone 38.9 34.7 33.2 26.1 28.1 30.6 26.9 29.4 31.4 31.0
Text-SFT 62.3 43.8 39.2 32.9 35.1 42.7 32.8 37.4 39.8 41.0
$M$-RL 63.8 51.7 45.8 35.3 38.0 44.4 39.1 42.2 44.7 46.9
OVR-7B Backbone 38.9 34.7 33.2 26.1 28.1 30.6 26.9 29.4 31.4 31.0
Text-SFT 66.2 46.7 38.7 33.8 33.7 37.9 32.5 38.1 40.7 42.9
$M$-RL 70.4 49.8 43.7 38.4 38.2 44.9 36.2 39.7 44.9 49.9
GLM-4.1V-9B Backbone 40.0 44.6 44.7 37.7 37.8 44.7 36.7 40.4 43.7 39.7
$M$-RL 69.2 65.8 59.9 52.3 50.2 60.8 50.0 57.0 61.4 62.8
VL-Rethinker-72B Backbone 52.9 41.6 41.0 36.2 35.5 42.2 37.5 38.4 41.6 40.0
$M$-RL 56.2 47.6 45.0 38.2 39.1 45.5 38.4 44.6 44.8 43.9
OpenAI 4O / O3 Backbone 49.1 40.7 39.1 35.2 32.9 40.7 33.6 39.0 40.8 38.7
Thinking 74.5 66.8 61.4 51.9 49.9 61.7 55.8 57.8 62.9 59.0
Gemini-2.5-Flash Backbone 79.9 69.2
Thinking 82.3 72.7
Claude-4-Sonnet Backbone 75.1 60.7 61.2 51.0 53.3 60.4 48.1 57.0 60.6 55.7
Thinking 84.1 65.6 64.4 51.0 55.0 65.6 51.5 61.0 67.0 57.0
Table 3: Downstream accuracy across diagram modifications with output consistency under these conditions. Text indicates performance from textual descriptions instead of diagrams, Raw denotes original human-generated diagrams, Base the newly rendered diagrams, and the other cases semantic-space modifications followed by rendering.
Downstream performance under diagram modifications.
Table˜3 reports the complete downstream evaluation results on MathLens, including all diagram modifications. These values underlie Figure˜4Figure˜6,  Figure˜9, and Figure˜10. Note that the error-type analysis in Figure˜14 relies on per-sample categorization and cannot be obtained directly from the aggregate scores presented here.
Model Variant Tri- Quad- Parallel Perpen- Collinear Co- Same len. Val. len. Same $\angle$ Val. $\angle$
-angle -rilateral -dicular -circular
VL-Rethinker-7B Backbone 69.1 52.3 49.0 78.2 59.5 65.6 57.4 86.6 50.0 68.1
$M$-RL 67.6 55.2 60.6 76.6 67.4 81.7 50.0 86.9 50.0 75.8
ShareVL-R1-7B Backbone 69.1 52.3 49.0 78.2 59.5 65.6 57.4 86.6 50.0 68.1
$M$-RL 71.4 57.7 62.5 78.2 66.0 88.9 57.4 86.4 69.2 72.4
R1-OneVision-7B Backbone 69.1 52.3 49.0 78.2 59.5 65.6 57.4 86.6 50.0 68.1
$M$-SFT 67.6 48.1 56.7 67.4 53.2 60.0 48.5 76.7 46.2 63.5
$M$-RL 71.4 54.4 59.6 72.2 61.1 59.4 55.9 79.1 61.5 66.9
Vision-R1-7B Backbone 69.1 52.3 49.0 78.2 59.5 65.6 57.4 86.6 50.0 68.1
$M$-SFT 65.8 49.4 43.3 73.5 64.8 60.0 45.6 78.6 61.5 73.8
$M$-RL 70.9 56.1 49.0 76.1 65.7 77.8 57.4 88.9 61.5 76.9
Revisual-R1-7B Backbone 69.1 52.3 49.0 78.2 59.5 65.6 57.4 86.6 50.0 68.1
Text-SFT 79.1 64.0 65.4 77.5 59.8 67.8 57.4 81.7 69.2 82.3
$M$-RL 84.3 72.0 65.4 82.8 68.0 77.2 60.3 88.1 84.6 87.8
OVR-7B Backbone 69.1 52.3 49.0 78.2 59.5 65.6 57.4 86.6 50.0 68.1
Text-SFT 83.5 61.5 79.8 82.2 61.0 69.4 57.4 83.2 69.2 86.0
$M$-RL 82.2 67.8 76.9 85.9 66.9 80.6 50.0 83.0 80.8 89.0
GLM-4.1V-9B Backbone 74.3 52.3 76.9 77.8 68.2 74.4 57.4 90.1 53.8 74.5
$M$-RL 82.2 66.5 88.5 88.8 66.2 91.7 82.4 95.3 92.3 88.0
VL-Rethinker-72B Backbone 66.7 49.8 76.9 80.4 66.8 94.4 67.6 94.5 69.2 87.4
$M$-RL 83.3 57.7 79.8 84.3 65.2 96.7 64.7 93.7 80.8 86.7
OpenAI 4O / O3 Backbone 81.9 60.3 78.8 80.3 69.5 95.0 67.6 93.4 80.8 87.5
Thinking 94.0 95.4 95.2 92.8 94.1 93.9 95.6 98.9 100.0 96.7
Gemini-2.5-Flash Backbone 92.7 90.0 97.1 92.8 89.5 96.1 95.6 97.5 88.5 92.1
Thinking 93.2 92.1 97.1 95.5 91.8 96.7 98.5 98.4 92.3 93.9
Claude-4-Sonnet Backbone 88.6 75.3 79.8 85.2 82.3 95.0 77.9 98.1 84.6 92.5
Thinking 93.5 93.7 85.6 88.8 83.1 89.4 83.8 98.5 88.5 93.7
Table 4: Perception probe accuracy by question types.
Perception probe results by question type.
Table˜4 reports the full benchmark scores for each perception probe question type, providing the numerical values underlying Figure˜13.
Figure 14: Distribution of error types per model. Most gains come from cases failing in both perception and reasoning. Integration errors become more prominent as other categories improve.
Error type analysis.
Figure˜14 show full error type analysis results for all models tested in this work.
E.4 MathLens-NS
Figure 15: Data samples from MathLens-NS. We curate problem instance from a prior dataset ((Fu et al., 2024)) and annotate the text description and perception-centric question-answers.
Figure 16: Data samples from MathLens-NS. We curate problem instance from a prior dataset ((Hao et al., 2025)) and annotate the text description and perception-centric question-answers.
Figure 17: Error type distribution across models on MathLens-NS. Most shifts are associated with perception-related cases.
Figure˜17 presents the error-type distribution for MathLens-NS. Consistent with the main experiment (Section˜3.4), most effects of multimodal reasoning finetuning concentrate on perception-related cases (Perception & Reasoning or Perception). However, these gains are less stable, reflecting that MathLens-NS spans broader domains than MathLens and often demands out-of-distribution generalization from the multimodal training sets. An exception is textual SFT models, which show substantial reductions in pure Reasoning errors. This indicates that, unlike in math geometry tasks, the diverse reasoning skills required for MathLens-NS are not well represented in the backbone (Qwen-2.5-VL). Finally, the higher fraction of Trivial correct cases arises from MathLens-NS ’s multiple-choice format, in contrast to the open-ended geometry subset MathLens.
Appendix F Qualitative Examples
Sample outputs on downstream geometry problems.
We compare the fine-tuned multimodal reasoners with their corresponding backbone MLLMs. Figure˜18 and Figure˜19 present cases where fine-tuning corrected an initially wrong answer, while Figure˜20 and Figure˜21 illustrate cases where the model produced incorrect answers both before and after fine-tuning.
Sample outputs on perception probes.
Figure˜22Figure˜23, and Figure˜24 show cases where textual SFT corrected initially wrong answers. Consistent with the quantitative results in Section˜3.3, the stronger cognitive patterns induced by textual SFT also promote improved perception.
Figure 18: Model responses for a downstream geometry problem in MathLens. The backbone model fails in perception, leading to an incorrect answer. The textual SFT model initially makes the same mistake but recovers the perceptual detail and produces the correct answer.
Figure 19: Model responses for a downstream geometry problem in MathLens. The backbone model exhibits a reasoning error, whereas the multimodal RL model provides the correct solution.
Figure 20: Model responses for a downstream geometry problem in MathLens. The multimodal RL model inherits the backbone model’s flawed reasoning, leading to perception errors and downstream failure.
Figure 21: Model responses for a downstream geometry problem in MathLens. The multimodal RL model corrects the perceptual error of the backbone model but ultimately fails to produce the correct answer due to a reasoning error.
Figure 22: Model responses for a perception probe in MathLens. The textual SFT model corrects the perceptual error present in the backbone model.
Figure 23: Model responses for a perception probe in MathLens. The textual SFT model corrects the perceptual error present in the backbone model.
Figure 24: Model responses for a perception probe in MathLens. The textual SFT model corrects the perceptual error present in the backbone model.